OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (5, 2); m = 3 y + 2 = 3(x - 5) y + 2 = 3x - 5 y - 2 = 3(x + 5) y - 2 = 3(x - 5)

OpenStudy (okra77):

just sub into the equation y-y1 =m(x-x1) so it will be y-2=3(x-5)

OpenStudy (anonymous):

Thank you cx